E9-20A (similar to) Question Help Space Corp. issued a $550,000, 8 percent mortgage on January 1, 2018, to purchase warehouses. Read the requirements Requirement 1. Complete the amortization schedule for Space Corp., assuming payments are made semiannually. (Round amounts to the nearest dollar.) Date Jan 1, 2018 Jun 30, 2018 Payment Interest Principal Loan Balance 550,000 $26,602 Requirements 1. Complete the amortization schedule for Space Corp., assuming payments are made semiannually. Round amounts to the nearest dollar. 2. Record the journal entries for (a) issuance of mortgage on January 1, 2018, and (b) the first semiannual payment on June 30, 2018. Print Done Enter any number in the edit fields and then click Check Answer
Answer #1

1)

Total payment = $26602

interest = 550000 × 8% × 6/12 = $22000

Principal payment = 26602 - 22000 = $4602

Loan balance = $550000 - $4602 = $545398

2)Journal entries

a) warehouse a/c dr. 550000

To mortgage loan a/c. 550000

b) loan a/c dr. 4602

Interest a/c dr. 22000

To. Bank a/c. 26602

For any query please comment and do give positive rating